Lead (IV) oxide is an important component of lead-acid car batteries. What volume of a 2.00mol/L solution of nitric acid is sufficient to react with 105g of lead (IV) oxide?


1
Expert's answer
2021-04-23T07:32:12-0400

2PbO2 + 4HNO3 → 2Pb(NO3)2 + 2H2O + O.


"Moles (PbO_2)= \frac{105}{239.2}=0.44moles"


"Moles (HNO_3)= \frac{0.44×4}{2}=0.88moles"


"Volume = \frac{0.88}{2.00}=0.44L"